WHT Checks

Hot Links


  1. First check the Operations Logbook for any restrictions that may be still in force. Turn on the DOME LIGHTS from the console and raise the control room blind. Avoid use of flood lights if possible. They are a big consumer and a source of heat.
  2. Make a visual inspection of the observing floor to check that the telescope is free to move. e.g. No ladders or steps are close to the telescope or test instruments connected.
  3. Push the LAMP TEST button on the console. Check that all used indicators are working (e.g. the ALARMS).
    Replace any defective lamps with new replacements (28v).

Preparing the telescope for checking

  1. Check that the MODE key on the engineering desk is set to ENG.
  2. Switch the BEARING OIL PUMP and G/B OIL on and wait for the lamps on the alarm panel to go out.
  3. At the bottom of the blue Marconi cabinet (Bay 6) containing the power amplifiers, switch on the following breakers:

If working at Cassegrain focus:

If working at Prime focus then switch on the breakers:

If working at either of the Nasmyth foci, the breakers for the Nasmyth rotators (Bay 4) need switching on:

Logging out from the WHT Observing System

The WHT observing system comprises of several computers. These being:

Shutting down the ICS and DAS computers

Logging out from the TCS

Starting the Telescope Control Systems (TCS) and checking the telescope

The TCS runs using VMS on a DEC ALPHA computer. Generic name: whttcs (lpas4)

  1. Double click on the OSADISPLAY2 'lpas4' icon and login as WHT_LOGIN.
  2. On the TCS start menu select 'START'. This window will disappear and two TCS windows will come up, the 'TCS DISPLAY' and the 'TCS_USER' window.
  3. If for some reason the absolute encoder is not working you need to zeroset on the target. This point can usually be skipped.
    In engineering mode move in AZ to 298 degrees, this is below the mark on the synchro position indicator. Make sure you are in the 'GREEN' zone. Type the command:

    USER> zeroset azimuth target
    Slowly move in positive direction until you hear the 'beep' of the TCS, detecting the zeroset.
  4. The TCS on startup defaults to the CASS as focal station and ISIS as instrument. It is important you choose the appropriate STATION, INSTRUMENT and AGSELECT:
    The station command chooses the appropriate rotator.
  5. Set the Autoguider switch as required for the night.
  6. On the Engineering console, turn the MODE key to COMP and press COMP RESET.
    n.b. Check that the meters showing the AZIMUTH and ALTITUDE motor currents are registering. One showing a +ve current, the other the same value, but in the -ve direction. These should be steady with no oscillations.
  7. Enter the coordinates of a position NW (1 hour less than the current Local Sidereal Time in RA and 40 degree Declination. Format for entering telescope co-ordinates is :
  8. USER> source name hh mm ss dd mm ss equinox


    name        No more than 20 characters
    hh mm ss    RA  hrs mins secs
    dd mm ss    DEC +/- degs mins secs
    equinox     Usually J2000 or A (apparent)

    Example: LST is at 21:30:15

    USER> source TEST 20 30 15 40 00 00 J2000
    USER> add
  9. Instruct the telescope to move to the demanded position with:
    USER> next
  10. To avoid loss of time you may tell the rotator to track on the current position:
    USER> rotator float
  11. The telescope should slew to the demanded position with the dome moving also.
  12. When TCS info display indicates that the telescope and the rotator are TRACKING, check the ENCODERS page by entering:
  13. USER> page enc
  14. Check that the altitude and rotator incremental (gear) and absolute encoders are in close agreement and updating and that the dome is tracking. The dome and telescope azimuth positions on the TCS info display should be similar within a few degrees. n.b. If the DS (or CWS) rotator does does NOT move? Has the DS or CWS breaker on the Nasmyth cabinet power control panel tripped or perhaps it has not been switched on?
  15. Entering:
    USER> page info
    will return the TCS Infodisplay window to the default page.
    n.b. If the dome does not move, it may have been overridden? On the console there is a bank of five push buttons labelled SELECT MANUAL CONTROLS. Check that the button DOME/SHUTTERS has not been set. If the dome still doesn't move, check that the keyswitch on the grey dome control cabinet (which rides with the dome) is in the REMOTE position. It may have been switched to LOCAL.
  16. When you have established that the telescope and rotators move and go into  ”TRACKING” carry out a servo test of the Alt, AZ and rotator which is current selected. At the TCS prompt type:
    USER> rms servo on
    and leave it running for 1 minute before typing:
    USER> rms servo off
    Record the RMS values in the LOG book.
  17. Before doing this, make a note of the current position of the focus. Move the focus a small amount then return it to it's previous value e.g. Focus is at 97.95 then enter say:
  18. USER> focus 98.25
  19. Check that it gets there then return it back.
  20. USER> focus 97.95
  21. At this point open the mirror petals. Go to the telescope and walk around the telescope fork. Check:
      - the dome is aligned &
      - all mirror petals are open.
  22. At this point you may decide to continue with testing the observing system leaving the telescope tracking. This has the advantage that taking exposures you can check that the telescope headers are recorded correctly.
  23. Once the headers are there you may park the telescope.

Completion of Telescope Checks

  1. Put the telescope back to zenith park by entering:
  2. USER> park zen
  3. To return the TCS to engineering mode, type:
  4. USER> eng
    n.b. You can also return the TCS to engineering mode by turning the MODE key to ENG. If this is done type ACK at the user prompt to cancel the alarm.
  5. Switch off the BEARING OIL PUMP and G/B OIL pump and press the ACCEPT button to cancel alarms as they sound (usually 3 times).
  6. Switch off the power amplifier circuit breakers.
  7. n.b. After a period of bad weather, the dome shutters should be exercised if ice/snow permits (check with mechanical section). Move the lower shutter down 50cm in micro-drive. Then open the upper shutter say 50 cm and close it again. Let the upper shutter reach the end stop. Finally close up the lower shutter. It isn't necessary to test dome shutter operation while the weather stays good, OSA's then exercise the dome shutters later opening the dome.

Starting the WHT observing system

The WHT observing system comprises as many DAS machines as cameras. Check the magnetic CCD board in the WHT control room, which gives up to date allocations of machines and CCD’s.

Use KDE (default) for WHTICSDISPLAY, AG DAS and TV DAS machines (see magnetic board)!

More info can be found in the:  Operations manual for UltraDAS.

  1. Check and update the configuration using http://taurus.roque.ing.iac.es:8081/Dewars/. The policy is to configure only what is used for the night. ACAM should always be configured if a CASS instrument is used. Do not configure test cameras.
  2. On the combined orange DAS window select the 1st DAS machines in use (see magnetic board).
  3. Enter : DAS> obssys (or (if instructed by software) 'obssys old')
  4. Now type: DAS> startobssys
    The DAS software will be loaded and the  DAS>  prompt returned when completed.
    n.b. The above procedure must be done in EACH orange DAS window for the detector(s) in use.
  5. In the WHTICS window enter: SYS> obssys (or (if instructed by software) 'obssys old')
  6. Now type: SYS> startobssys
    During start-up, various windows will be generated. The important ones being:
    • Talker
    • Observing log
    • combined CCD status monitor
    • Who goes there (observer’s info)
  7. When the SYS> prompt is returned. The observing system is ready to rock n'roll!
  8. To avoid screen clutter, minimise the windows that are not being used.

Taking a test bias image

  1. Restore CCD settings of all cameras to saved values: setccd  <camera>      (e.g. SYS> setccd red)
  2. Before taking images check the readout speed is set to 'slow':
    SYS> rspeed channel slow   (where channel = red; blue; acam; pfip; oasis; ingrid; wyffos; integral; etc)
  3. Ensure nobody is in the dome without a torch and switch all dome lights off.
  4. To take an image, enter at the prompt:
    SYS> bias channel
  5. Wait until the CCD status monitor indicates that the CCD has been read out and has been 100% processed. Make sure the image gets written (with instrument headers and TCS header information when tracking) to the Observing Log.

Displaying the image

  1. If not already done, login to the WHT Data Reduction PC (whtdrpc1) as:  whtguest
  2. Open a teminal and launch the command 'iraf'. This script starts up all you need!
  3. Within IRAF, change to the obsdata directory where the images are being stored. Note, a new ’date’ directory is generated at 10:00 UTC daily. e.g.

    cd /obsdata/whta/20100521

    The combined CCD status monitor window contains information on which disc the images are being stored.
    Each image (or image pair if using the Mosaic camera) is contained in a single FITS file with: rnnnnnnn  being the last run number.

  4. To display the image, enter at the IRAF prompt:

    display r1075678[1] 1 fill+

Examining the image

  1. At the IRAF prompt, enter:  imexam
  2. Move the mouse into the area of the displayed frame and left click. Move the mouse around the image area, typing "m".  This will produce statistics for a box containing 25 pixels centred on the mouse cursor. Check that the noise(e) levels correspond to the CCD data held in:
    http://www.ing.iac.es/ Engineering/engweb6a.htm
  3. To exit imexam, leave the mouse cursor over the image area and type:  q

Checking autoguider & TV systems

  1. Use KDE (default) for WHTICSDISPLAY, AG DAS and TV DAS machines (see magnetic board)!
  2. Enter: DAS> obssys followed by
  3. DAS> startobssys
  4. Start the autoguider and the TV using the ‘startag <autoguider_name> ’ script. For example to start the ag6 autoguider, then type: startag AG6

    During start-up, various windows will be generated. The important ones being:

    • Talker
    • Autoguider GUI
    • DS9

  5. Take a field exposure and check light is registered.

  6. The Artificial Guide Star Test at Cass (please perform this test!)

    A more in-depth test for the autoguider is using the artificial guide star. It should be noted that this facility is only available for the CASS autoguider, here it allows the autoguider error correcting loop to the TCS to be closed.

    n.b. The TCS must be running with the telescope TRACKING to close the loop.

    Deploy the opaque filter with the pin hole:

    autofilt ag-opaque

    To see the artificial star, some stray light is required. Turning on just the FORK lights on the telescope should suffice. If more counts are required, move in the comparison lamp mirror and turn on the tun(white) lamp and take an exposure. The ICS commands being:

    complamps w
    DAS> field (from the autoguider window)

    After reading out, the artificial guide star will be clearly visible with a guide box and cross positioned over it.

    To initiate guiding, enter:

    DAS> guide on (from the autoguider window)

    With the TCS running and the telescope tracking, to close the guide loop enter:
    USER> autoguide on

    Check the TCS Infodisplay to see that autoguiding has been enabled.

    To restore the autoguider for normal operations after carrying out this test. Enter the commands:
    USER> autoguide off
    DAS > guide off
    complamps off (if on)
    autofilt ag-clear

Instrument Checks

When checking out the instruments, use commands that will NOT move mechanisms that have been setup by the observers. This particularly applies to COLLIMATORS and GRATINGS. Moving filters, autoguider probes etc. is the safest option.


  1. Cass A&G Box

    prag 15000 50000 (moves the Cass autoguider probe)
    compnd 1 (put in a neutral density filter)
    compnd 0 (remove it)


  2. ISIS

    Take a note of the dekker position (usually '8'), move the dekker slide and put it back. E.g.:

    dekker 7 (move dekker slide to postion 7)
    dekker 8 (return to original position)

    hwin (move in the half wave plate, ISISP check)
    hwout (remove it)

    Take ISIS arc images

    1. Move in the comparison mirror and switch on both arc lamps:

      agcomp (moves the comparison mirror in)
      complamps CuNe+CuAr (switches both arc lamps on)

      We assume the grating angles are set to practical values. NOTE: Do not move COLLIMATORS and GRATINGS during an observing run.

      Check the slit width but don't change it. With a very wide slit, arc lines will come out broad.


    2. At the SYS> prompt, take an exposure (preferably with dome lights out) in each arm deemed to be available for the night e.g.

      arc blue 20 &
      arc red 10 &

      The combined CCD status window will show the state of progress as the exposure is read out and is processed. Make sure the image gets written (with instrument headers and TCS header information when tracking) to the Observing Log.


    3. Check with IRAF displaying and examining the image that arc lines are clearly visible.


  3. ACAM

    ACAM daily checks


  4. PFIP

    agprobe 50000 10000 (moves the PF autoguider probe)
    pfilter 3 (moves main filter wheel. Range is 1 to 7)
    pfip_status (displays the status of the system)

    To display both chips you may window, including both chips:
    window pfip 1 "[1:4300,1:4200]"

    With the TCS running, move the telescope to different altitudes and check that the ADC elements move. The status associated with the position of either of the ADC elements should be blue as the ADC changes position.

    On the TO terminal, start up a telnet session on taurus as whtobs. Next type obssys and select the same option as was selected when the observing system was selected on taurus.

    Execute a pfip_status command from this terminal to ensure that communications are working correctly. A display of the status of the instrument should be displayed.

    Take PFIP dome flat image

    1. Open the mirror petals and switch on the low power top ring flat field lamp.
      (Set the intensity to 3/4 (eng. console); with PFIP there is no need for any of the three 500W lamps).


    2. Take a flat field exposure

      flat pfip 5 &

      The combined CCD status window will show the state of progress as the exposure is read out and is processed. Make sure the image gets written (with instrument headers and TCS header information when tracking) to the Observing Log.


    3. Check with IRAF displaying and examining the image that a uniform light level is registered.


Visitor instruments on CASS: Instrument checks are performed by the visiting team. These should be carried out well before ING day time staff leaves. The CAGB and ACAM however must be tested! ACAM is usually cold and is by default required for override programs.


For checking our other instruments refer to the following documents. These additional docs give the full check-out procedure for these instruments:

ACAM daily checks

LIRIS daily checks

AF2 daily checks

BLT daily checks

NAOMI daily checks

INGRID daily checks

OSCA: This is rarely used. With the telescope tracking check on the OSCA GUI that the Lyot Angle is updating. Do ask the instrument specialist for advice.

OASIS daily checks

INTEGRAL daily checks


Filling the WHT Cryostats

Most detectors are now filled by LN2 auto-fill systems. Although some detectors may not be required for observing, they may need to still be kept cold. If you have any doubt, check the CCD status board in the middle of the console to see what cryostats need to be maintained cold or seek advice from the detector specialist.


E.g. before moving the telescope to AP3 ensure the CASS rotator position is at:

With PFIP move the rotator to:
USER> ROT MOUNT -75   before lowering the telescope!

Wear safety helmet (with visor), apron and gloves when filling cryostats manually or checking LN2 auto-fill!

  1. On the dewar top, check that the SILVER valve (gas vent) is closed! Vessels must be left pressurised.
  2. Use the indicator of the dewar to check the dewar contains enough LN2 (>25%). If it is lower tell the operations leader.
  3. Wheel the dewar carefully into position.
  4. Connect the dewar to earth using the established earthing points.
  5. FULLY insert the filler tube into the cryostat.
  6. On the dewar top, OPEN the corresponding BLUE decant valve; unless the pressure is considerably below 1 Bar, there is no need to open the GREEN pressure build valve.
  7. Wait until liquid nitrogen is seen to be coming out of the cryostat.
  8. CLOSE the green and blue valves.
  9. Wait for the filler tube to warm up (until it is not rigid anymore) and remove it from the cryostat.
  10. Move the dewar back to the park position.
  11. Note down in the WHT Logbook what time the cryostat was filled and the initials of your name.

Before leaving the WHT

Back in the control room. Turn OFF the dome lights and don't forget to tick off the checks as completed and to sign off in the logbook.

Document History

Rev 3.0 July 27th 1992 CRB
Rev 5.0 Feb 12th 1994 CRB
Rev 6.0 Feb 19th 1999 CRB
Rev 7.0 Sep 16th 1999 EJM  CRB's edition re-written in HTML format
Rev 8.0 Jun  14th 2001 EJM  Completely revised to cover UltraDAS
Rev 8.1 Jun  27th 2001 EJM  Minor changes made
Rev 8.2 Oct 15th 2001 EJM/SHEILA  Config file check added
Rev 8.3 Dec 12th 2001 EJM   Aux port is now on UDAS
Rev 8.4 July 16th 2002 EJM   Machine names removed. Generic names only now
Rev 8.5 Sep 12th 2002 EJM   Logging off WHT Observing System added
Rev 8.6 Sep 17th 2002 EJM   References to UES removed
Rev 8.7 Oct 9th 2002 EJM   PNS channel selection added
Rev 8.8 Dec 18th 2003 CB/FJG Added details relating to new UNIX PFIP control system & TCS start
Rev 8.9 Jun 28th 2004 KMD/RAB Added details of SDSU autoguiders and TVs. Westinghouse info removed.
Rev 9.0 Dec 9th 2004. KMD added RMS servo Checks and removed cryocam instructions.
Rev 10.0 Oct 27th 2007. JR. New structured to the make checks more efficient, added links to all WHT instruments.

Rev 10.1 Apr 25th 2008. JR. added saveccd & setccd.

Rev 10.2 Jun 19th 2008. JR. ISIS checks: don't alter slit width.

Rev 10.3 Oct 30th 2008. JR. ISIS checks: updated dekker movement.

Rev 10.4 Nov 19th 2008. JR. logout as whtobs from whticsdisplay.

Rev 10.5 Jun 2nd 2009. JR. OSCA: check Lyot Angle.

Rev 10.6 Jun 9th 2009. JR. added ACAM daily checks.

Rev 10.7 May 21st 2010. JR. added 'DEWARS' configuration utility.

Rev 11 June 6th 2014. JR. General revision.

Rev 12 June 26th 2015. JR. Emphasis made on header information.

Click here for the WHT home page.

THIS FILE: http://www.ing.iac.es/~eng/ops/wht/wht_checks.html